Laser and Locking troubleshooting

From Hall A Wiki
Jump to: navigation, search

The laser system is functioning in a non-optimal mode due to issues with the fiber amplifier. The normal laser setup and locking procedure is:

  • Turn on seed laser
  • Turn on fiber amplifier at 3 W
  • Optimize PPLN temperature
  • Lock cavity